Question 5 Use Appendix Table Ш to determine to 5 decimal places the following probabilities for the standard normal random variable Z: (a) P(Z< 1.29) = (b) P(Z < 2.8) = (c) PlZ > 1.45) = (a) PIZ> 2.15) (e) P(-2.34 < Z < 1.76) =
Answer #1

Solution :

Using standard normla table,

a ) P(Z < 1.29 )  

P(Z < 1.29 ) = 0.9015

b )  P(Z < 2.8 )  

P(Z < 2.8 ) = 0.9974

c )  P(Z > 1.45 )  

= 1 - P(Z < 1.45 )  

= 1 - 0.9265

= 0.0735

P(Z > 1.45 ) = 0.0735

d )  P(Z > - 2.15 )  

= 1 - P(Z < - 2.15 )  

= 1 - 0.0158

= 0.9842

P(Z > - 2.15 ) = 0.9842

e )  P(- 2.34 < Z < 1.76 )  

= P(Z < 1.76 ) - P ( Z < -2.34 )  

= 0.9608 - 0.0096

= 0.9512

P(- 2.34 < Z < 1.76 ) = 0.9512